# From System 1 to System 2: A Survey of Reasoning Large Language Models

### v2

@@ 5.1 Efficient Reasoning LLMs
While reasoning LLMs excel at solving complex problems via extended inference, their reliance on long autoregressive reasoning within large-scale architectures presents significant efficiency challenges. For example, many problems on platforms like Codeforces require over 10,000 tokens of reasoning, resulting in high latency. As noted in [122], even when a reasoning LLM identifies the correct solution early, it often spends considerable time verifying its reasoning. Recent reports, such as Deepseek-R1 [31], suggest that self-improvement via RL is more effective in larger models, while smaller-scale large language models (SLMs) (e.g., 3B and 7B models as explored by [123] and [239, 222]) struggle to match performance in slow-thinking reasoning tasks.
@@ 5.5 Multilingual Reasoning LLMs
5.6 Multimodal Reasoning LLMs
Extending slow-thinking reasoning capabilities from text-based domains to multimodal contexts remains a significant challenge, especially in tasks requiring fine-grained perception [116, 295, 296]. While approaches like Virgo [297] have attempted to distill text-based slow-thinking reasoning into multimodal LLMs, their performance improvements in tasks such as MathVision [262], which demand detailed visual understanding, have been marginal.
Key research directions include developing hierarchical reasoning LLMs that enable fine-grained cross-modal understanding and generation, tailored to the unique characteristics of modalities such as audio, video, and 3D data.
@@ 5.7 Safe Reasoning LLMs
The rapid development of reasoning LLMs like OpenAI-o1 [29] and DeepSeek-R1 [31] has led to the rise of superintelligent models capable of continuous self-evolution. However, this progress brings challenges in safety and control [298, 299, 300]. RL, a key training method, introduces risks such as reward hacking, generalization failures, and language mixing, which can lead to harmful outcomes. Ensuring the safety of such systems like DeepSeek-R1 is urgent. While RL enhances reasoning, its uncontrollable nature raises concerns about safely guiding these models. SFT addresses some issues but is not a complete solution. A hybrid approach combining RL and SFT is needed to reduce harmful outputs while maintaining model effectiveness [301].

### v3

@@ 6.1 Efficient Reasoning LLMs
While reasoning LLMs excel at solving complex problems via extended inference, their reliance on long autoregressive reasoning within large-scale architectures presents significant efficiency challenges [493, 494]. For example, many problems on platforms like Codeforces require over 10,000 tokens of reasoning, resulting in high latency. As noted in [124], even when a reasoning LLM identifies the correct solution early, it often spends considerable time verifying its reasoning. Recent reports, such as Deepseek-R1 [31], suggest that self-improvement via RL is more effective in larger models, while smaller-scale large language models (SLMs) (e.g., 3B and 7B models as explored by [125] and [286, 262]) struggle to match performance in slow-thinking reasoning tasks.
@@ 6.6 Safe Reasoning LLMs
The rapid development of reasoning LLMs like OpenAI-o1 [29] and DeepSeek-R1 [31] has led to the rise of superintelligent models capable of continuous self-evolution. However, this progress brings challenges in safety and control [508, 509, 510]. RL, a key training method, introduces risks such as reward hacking, generalization failures, and language mixing, which can lead to harmful outcomes. Ensuring the safety of such systems like DeepSeek-R1 is urgent. While RL enhances reasoning, its uncontrollable nature raises concerns about safely guiding these models. SFT addresses some issues but is not a complete solution. A hybrid approach combining RL and SFT is needed to reduce harmful outputs while maintaining model effectiveness [511, 512].